# Random local access for sampling $k$-SAT solutions

### v1

@@ 4.2 Efficiency
Here in the last inequality, we used the fact that $\theta>0.4$ implied by 3 and the assumption that $d\leq 2^{\alpha k/4}$. Setting $b=kd\log^{2}n$, we get that $\mathbb{P}[\text{$\Phi_{v}^{\sigma}$ has $b$ clauses}]\leq 0.9^{\log^{2}n}\leq n^{-0.1\log n}.$
With Proposition 4.4, we show that Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation for every marked variable.
Proposition 4.5.
Suppose $d\leq 2^{\alpha k/4}$. Let $\Phi$ be a $(k,d)$-formula satisfying Lemma 2.11. Suppose $(\mathsf{IsMarked}(v))_{v\in V}$ gives a valid marking $\mathcal{M}$. Then for all $v\in\mathcal{M}$, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute.
Proof.
Suppose we run Algorithm 1 on a marked variable $v$. Then we have the following breakdown of the complexity of Algorithm 1:
  1. 1.
    Computing $\mathsf{IsMarked}(v)$. By Theorem 2.12, this takes $\mathrm{polylog}(n)$ time in expectation.
  2. 2.
    In $\text{MarginSample}(\Phi,\emptyset,v)$, we first sample $r_{v}\in[0,1]$ uniformly at random.
    • If $r_{v}\leq 2\theta$, then we choose $\sigma(v)\in\{0,1\}$ uniformly at random, which takes constant time.
    • If $r_{v}>2\theta$, then we call $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$. Let $s$ denote the number of clauses in $\Phi_{v}^{\sigma}=\mathsf{Conn}(\Phi,\sigma,v)$ in line 9 of $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$. By Proposition 4.4, we know that
      $\mathbb{E}[s]\leq n^{-0.1\log n}|\mathcal{C}|+(1-n^{-0.1\log n})kd\log^{2}n\leq n^{-0.1\log n}dn+kd\log^{2}n\leq 2kd\log^{2}n.$
      Given that $\Phi_{v}^{\sigma}$ has $s$ clauses, we know that there are at most $O(s)$ further iterations of $\mathsf{MarginOverflow}$ within $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$ as $\Phi$ has constant maximum degree. Each iteration of $\mathsf{MarginOverflow}$ involves computing $\mathsf{IsMarked}(w)$ for a constant number of variables $w$ (plus a constant number of random samplings), so by Theorem 2.12, it takes $\mathrm{polylog}(n)$ time in expectation. Summing up, we know that computing $\Phi_{v}^{\sigma}$ takes $\mathrm{polylog}(n)\cdot s$ time in expectation. Finally, by Theorem 4.3, we know that it takes $O(s)$ time in expectation to compute $\mathsf{UniformSample}(\Phi_{v}^{\sigma})$. This altogether tells us that $\mathsf{MarginOverflow}(\Phi,\emptyset_{v\leftarrow\star},v)$ has expected complexity at most $\mathrm{polylog}(n)$.
We now show that Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation for every unmarked variable as well.
Proposition 4.6.
Suppose $d\leq 2^{\alpha k/4}$. Let $\Phi$ be a $(k,d)$-formula satisfying Lemma 2.11. Suppose $(\mathsf{IsMarked}(\Phi,v))_{v\in V}$ gives a valid marking $\mathcal{M}$. Then, for all $v\in V\setminus\mathcal{M}$, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute.
Proof.
Let $\sigma\in\{0,1\}^{\mathcal{M}}$ denote the partial assignment Algorithm 1 would output on $\mathcal{M}$, were we to compute $\sigma_{v}$ for every $v\in\mathcal{M}$. By Theorem 4.1, we know that $\sigma$ follows the marginal distribution $\mu_{\mathcal{M}}$. Moreover, we know from Lemma 2.11 that with probability at least $1-\delta/n^{2}$, $\sigma$ $\delta$-shatters $\Phi$, which means that every connected component in $\Phi^{\sigma}$ has size $\lesssim dk\log(n/\delta)$. We will not actually construct the entire partial assignment $\sigma$, but will locally use this shattering property as described below.
For every $v\in V\setminus\mathcal{M}$, let $\Phi^{\sigma}_{v}$ denote the connected component in $\Phi^{\sigma}$ containing $v$. Note that $\Phi^{\sigma}_{v}=\mathsf{Conn}(\Phi,\emptyset,v)$ and thus can be locally computed without extending $\sigma$ to a full assignment.
Suppose $\Phi^{\sigma}_{v}$ has $s$ variables. By Theorem 2.12 and Proposition 4.5, both $\mathsf{IsMarked}(\cdot)$ and $\mathsf{MarginSample}(\Phi,\emptyset,w)$ (for any $w\in\mathcal{M}$) have expected $\mathrm{polylog}(n)$ complexity. Since we execute each of these two operations $O(s)$ times when we compute $\Phi_{v}=\mathsf{Conn}(\Phi,\emptyset,v)$, we know that $\mathsf{Conn}(\Phi,\emptyset,v)$ has expected $s\cdot\mathrm{polylog}(n)$ complexity. It remains to sample a uniformly random solution using $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$. By Theorem 4.3, we know that $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$ has expected $O(s)$ complexity.
Since $s\lesssim dk\log^{2}n$ with probability at least $1-\delta/n^{2}$, we get that in expectation,
  • $\mathsf{Conn}(\Phi,\sigma,v)$ has time complexity at most
    $(1-\delta/n^{2})\cdot dk\log^{2}n\cdot\mathrm{polylog}(n)+\delta/n^{2}\cdot n\cdot\mathrm{polylog}(n)=\mathrm{polylog}(n);$
  • $\mathsf{UniformSample}(\Phi^{\sigma}_{v})$ has complexity at most
    $(1-\delta/n^{2})\cdot O(dk\log^{2}n)+\delta/n^{2}\cdot O(n)=\mathrm{polylog}(n).$
Therefore, Algorithm 1 takes $\mathrm{polylog}(n)$ time in expectation to compute. ∎
4.3 Completing the proof
Having gathered all the ingredients, we can now prove the main theorem that is the formal version of 1.1.
Theorem 4.7.
Suppose $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula satisfying (1). Let $\mu$ be the uniform distribution over satisfying assignments to $\Phi$, with marginal distributions $\mu_{v}$ for $v\in V$. Then for all $c>0$, there is a $(\mathrm{polylog}n,n^{-c})$-random local access algorithm $\mathcal{A}$ for sampling the variable assignment of $v\in V$ as $\widehat{\mu}_{v}$, such that
Proof.
Suppose $\Phi=(V,\mathcal{C})$ is a $(k,d)$-formula satisfying (1). Fix constant $c>0$.
We first choose the parameters $\alpha,\beta_{1},\beta_{2}$ that satisfy Theorem 2.12. Let
$\alpha=1/75,\qquad\beta_{1}=0.778,\qquad\beta_{2}=0.96.$
One can verify that $1/2<\beta_{1}<\beta_{2}<1-\alpha$ and with these parameters, conditions in Theorem 2.12 are satisfied. Thus, by Theorem 2.12, there exists a $\mathrm{polylog}(n)$ oblivious local computation algorithm $\mathsf{IsMarked}(\cdot)$ that with probability at least $1-1/n$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$.
Suppose $\mathsf{IsMarked}(\cdot)$ gives a consistent $\alpha$-marking $\mathcal{M}\subseteq V$. Since $d\lesssim 2^{k/400}$, condition 4 is satisfied by parameters $\alpha,d,k$. By Proposition 4.2 and Theorem 4.3, we know that for all variable $v\in V$, Algorithm 1 on $v$ returns some $\widehat{\mu}_{v}$ within total variation distance $n^{-c}$ of $\mu_{v}$.
Since $d\lesssim 2^{k/400}$, our choice of $\alpha$ gives the condition $d\leq 2^{\alpha k/4}$ required by Propositions 4.5 and 4.6. These two propositions tell that Algorithm 1 on every $v\in V$ takes $\mathrm{polylog}(n)$ time in expectation.
By construction, Algorithm 1 is memory-less as it samples all necessary variable assignments in order to compute the assignment of a queried variable $v$. Furthermore, Algorithm 1 queried on different variables $v\in V$ collectively returns an assignment $\sigma\in\Omega$ that has $d_{\operatorname{\mathsf{TV}}}(\textsf{Law}(\sigma),\mu)<n^{-(c-1)}$ by construction. Since this holds for any $c>0$ constant, we obtain the desired result. ∎
5 Concluding remarks
With more involved refinements and optimizations of the arguments in this work, the density constraint $d\lesssim 2^{k/400}$ of 1.1 can be substantially improved (to something closer to $d\lesssim 2^{k/50}$). We omit these additional calculations in favor of expositional clarity to highlight our main result: random local access models exist for arbitrary bounded degree $k$-CNFs at exponential clause density. Furthermore, these arguments can also be adapted (in a similar fashion to [11, 17, 5]) to obtain a random local access models for random $k$-CNFs in a comparable density regime.
Nonetheless, the limit of the approaches in this work would still fall well short of obtaining random local access for, e.g., approximately $d\lesssim 2^{k/5}$, the largest density at which we currently know how to efficiently sample solutions to an arbitrary bounded degree $k$-CNF in nearly-linear time ([16]). This is because of our reliance on a query-oblivious LCA to construct a local marking and our application of sampling results to a correspondingly reduced CNF.
The approach we take in this work is only one of many possible schema to reduce from existing classical, parallel, and/or distributed algorithms to more local algorithms. Our approach involved using ideas and techniques from a variety of previous works (including [1, 16, 21]), many of which were in the non-local setting, and adapting and combining them in a novel way to obtain a sublinear sampler. In [4], the authors adapted a parallel Glauber dynamics algorithm for sampling colorings to obtain random local access to random proper $q$-colorings. We conjecture that one could, using similar ideas to [4], adapt parallel Glauber dynamics or algorithmic correlation decay (as studied in [9, 20]) in conjunction with our local marking scheme to obtain different sampling algorithms. We expect such algorithms would also provide random local access to satisfying assignments of a $k$-CNF at comparable clause density to this work.
Appendix A Proof of Theorem 2.12
@@ 5 Concluding remarks
Given a $k$-uniform hypergraph $H$ and constants $10<\alpha<\beta_{1}<\beta_{2}<$ 1, we describe an algorithm $\mathcal{A}$ that finds a 2-coloring of $H$, such that every edge in $H$ has at least $\alpha k$ vertices in each color. Taking $H=H_{\Phi}$ (as in 2.1), we see that finding an $\alpha$-marking for $\Phi$ is equivalent to finding such a 2-coloring for $H_{\Phi}$. The algorithm $\mathcal{A}$ runs in three phases.
@@
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k},$
$\delta\mapsto(\beta_{2}-\delta)-h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{(1-h(\frac{\alpha}{1-\beta_{2}}))(1-\beta_{2})k}.$
@@
Lemma A.1.
With probability at least $1-\frac{1}{n}$, after Phase 1, all connected components in $H$ have size at most $c_{1}\log n$.
@@
Let $W$ be a 3-tree on $w$ vertices, and $p=4kd^{2}2^{(h(1-\beta_{1})-\beta_{1})k}$. Then
@@
which is at most $1/n$ when $w=c_{1}\log n$. By Markov’s inequality, with probability at least $1-1/n$, there is no 3-tree of size larger than $c_{1}\log n$. Since every size-$s$ connected component in $H$ contains a 3-tree of size at least $s/(dk)^{3}$, we get that with probability at least $1-1/n$, there is no connected component of size larger than than $c_{1}(dk)^{3}\log n$ in $H$. ∎
Lemma A.2.
Let $H_{1}$ be a connected component of $H$ after Phase 1. Suppose $H_{1}$ has size at most $c_{1}\log n$. Then with probability at least $1-\frac{1}{\log^{2}n}$, after Phase 2, all connected components in $H_{1}$ have size at most $c_{2}\log\log n$.
@@
The argument in similar to the proof of Lemma A.1. Let $e$ be any edge, and $v$ be any vertex in $\widetilde{H}_{1}$. Observe that
@@
$\mathbb{P}[\text{$v$ becomes troubled-2 in Phase 2}]\leq 2d\binom{(1-\delta)k}{(\beta_{2}-\delta)k}2^{-(\beta_{2}-\delta)k},$
@@
$\qquad\leq 4kd^{2}2^{h(\frac{\beta_{2}-\delta}{1-\delta})(1-\delta)k-(\beta_{2}-\delta)k}\leq 4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}.$
Let $\widetilde{p}=4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}$. Thus, the expected number of 3-trees on $w$ vertices after Phase 1 is at most
$N4^{w}(dk)^{3w}\widetilde{p}^{w}$ $=N(4d^{3}k^{3}\cdot 4kd^{2}2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w}$
$=N(16k^{4}d^{5}\cdot 2^{h(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}})(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w},$
which is at most $1/N^{2}$ when $w=c_{2}\log N$. By Markov’s inequality, with probability at least $1-1/N^{2}$, there is no 3-tree of size larger than $c_{2}\log N$. This implies that with probability at least $1-1/\log^{2}n$, there is no connected component of size larger than than $c_{2}(dk)^{3}\log\log n$ in $H$. ∎
Lemma A.3.
@@
$\mathbb{P}[A_{e}]\leq 2\binom{(1-\beta_{2})k}{\leq\alpha k}2^{-(1-\beta_{2})k}\leq 2^{(h(\frac{\alpha}{1-\beta_{2}})-1)(1-\beta_{2})k+1}.$
@@
$e\cdot 2^{(h(\frac{\alpha}{1-\beta_{2}})-1)(1-\beta_{2})k+1}\cdot(kd+1)<1,$
@@
  • For every $w\in V$ that shares a clause with $v$, if $a_{w}<a_{x}$, then we add $w$ to $\mathcal{T}_{v}$ as a child of $v$.
@@
Lemma A.4 ([1, Theorem 3.1]).
For every $k,d\geq 1$, there exists $c_{3}>0$ such that for all sufficiently large $n$, $\mathbb{P}[|\mathcal{T}_{v}|>c_{3}\log^{kd+1}(n)]<1/n^{2}$.
In Phase 1, whether $v$ is colored or troubled-1 completely depends on its descendants in $\mathcal{T}_{v}$. Thus, to compute the assignment of $v$ in Phase 1, it suffices to run Phase 1 on all vertices in $\mathcal{T}_{v}$. By Lemma A.4, with probability $1-1/n^{2}$, this step takes $\mathrm{polylog}(n)$ time.
If $v$ gets colored in Phase 1, then we return the color directly. Otherwise, if $v$ becomes troubled-1 after Phase 1, then we compute the connected component of $H$ containing $v$ after Phase 1. Recall from Lemma A.1 that with probability $1-1/n$, all connected components after Phase 1 has size at most $c_{1}\log n$, so this step takes another $\mathrm{polylog}(n)$ time.
Let $H_{v}$ denote this connected component. By Lemma A.2, if we perform $O(\log n/\log\log n)$ parallel instances of Phase 2 on $H_{v}$, then with probability $1-1/n^{2}$, in one of them $H_{v}$ is further reduced to connected components each of size $c_{2}\log\log n$. We then perform Phase 3 to find a valid coloring on the component that contains $v$, which by Lemma A.3 is guaranteed to succeed and finishes in $2^{O(\log\log n)}=\mathrm{polylog}(n)$ time.

### v2

@@ Appendix B Efficiency of Algorithm 6
Here in the last inequality, we used the fact that $\theta>0.4$ and the assumption that $d\leq 2^{\alpha k/4}$. Setting $b=kd\log^{2}n$, we get that $\mathbb{P}[\text{$\Phi_{v}^{\tau}$ has $kd\log^{2}n$ clauses}]\leq 0.9^{\log^{2}n}\leq n^{-0.1\log n}.$
Observe that in $\mathsf{Conn}(\cdot)$, every clause is added to $\mathcal{S}$ because its variable set intersects with that of some unsatisfied clause in $\Phi_{v}^{\tau}$. Thus by Proposition C.1, we know that with probability $1-n^{-0.1\log n}$ the number of clauses ever added to $\mathcal{S}$ is at most $\Delta kd\log^{2}n\leq k^{2}d^{2}\log^{2}n$. In this case, Line 7 of Algorithm 7 is executed at most $k^{2}d^{2}\log^{2}n\cdot k=k^{3}d^{2}\log^{2}n$ times. Combined with Lemma 4.7, we get that the expected cost of $\mathsf{Conn}$ is at most
$(k^{3}d^{2}\log^{2}n+n^{-0.1\log n}\cdot n)\cdot O(k^{17}d^{10}\log^{2}n/\alpha)=\mathrm{poly}\log(n).$
Thus, Theorem 3.3 is proved.
Appendix D Proof of Theorem 2.10
@@
Let $c>0$ be any constant. Given a $k$-uniform hypergraph $H$ and constants $10<\alpha<\beta_{1}<\beta_{2}<$ 1, we describe a local computation algorithm $\mathcal{A}$ that with probability $1-n^{-c}$ finds a 2-coloring of $H$, such that every edge in $H$ has at least $\alpha k$ vertices in each color. Taking $H=H_{\Phi}$ (as in 2.2), we see that finding an $\alpha$-marking for $\Phi$ is equivalent to finding such a 2-coloring for $H_{\Phi}$. The algorithm $\mathcal{A}$ runs in three phases.
@@
$16k^{4}d^{5}\leq 2^{(\beta_{2}-\beta_{1})k-h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k},$
$\delta\mapsto(\beta_{2}-\delta)-h\left(\frac{\beta_{2}-\delta}{1-\delta}\right)(1-\delta)\text{ is decreasing on $0\leq\delta\leq\beta_{1}$,}$
$2e(kd+1)<2^{\left(1-h\left(\frac{\alpha}{1-\beta_{2}}\right)\right)(1-\beta_{2})k}.$
@@
Lemma D.1.
There exists $c_{1}>0$ sufficiently large such that with probability at least $1-n^{-2c}$, after Phase 1, all connected components in $H$ have size at most $c_{1}\log n$.
@@
Observe that every vertex that remains after Phase 1 lies in some dangerous-1 or unsafe-1 edge. Let $W$ be a 3-tree on $w$ vertices, and $p=4kd^{2}2^{(h(1-\beta_{1})-\beta_{1})k}$. Since no two vertices in $W$ lie in edges that intersect, we have
@@
which is at most $n^{-2c}$ when $w=c_{1}\log n$ with $c_{1}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-n^{-2c}$, there is no 3-tree of size larger than $c_{1}\log n$. Since every size-$s$ connected component in $H$ contains a 3-tree of size at least $s/(dk)^{3}$, we get that with probability at least $1-n^{-2c}$, there is no connected component of size larger than than $c_{1}(dk)^{3}\log n$ in $H$. ∎
Lemma D.2.
Let $H_{1}$ be a connected component of $H$ after Phase 1. Suppose $H_{1}$ has size $N\leq c_{1}\log n$ with $c_{1}>1$. Then there exists $c_{2}>0$ sufficiently large such that with probability at least $1-1/\log^{2}n$, after Phase 2, all connected components in $H_{1}$ have size at most $c_{2}\log\log n$.
@@
The argument in similar to the proof of Lemma D.1. Let $e$ be any edge, and $v$ be any vertex in $\widetilde{H}_{1}$. Observe that
@@
$\qquad\leq 4kd^{2}2^{h\left(\frac{\beta_{2}-\delta}{1-\delta}\right)(1-\delta)k-(\beta_{2}-\delta)k}\leq 4kd^{2}2^{h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}.$
Let $\widetilde{p}=4kd^{2}2^{h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k-(\beta_{2}-\beta_{1})k}$. Thus, the expected number of 3-trees on $w$ vertices after Phase 1 is at most
$N4^{w}(dk)^{3w}\widetilde{p}^{w}$ $=N(4d^{3}k^{3}\cdot 4kd^{2}2^{h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w}$
$=N(16k^{4}d^{5}\cdot 2^{h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k-(\beta_{2}-\beta_{1})k})^{w},$
which is at most $N^{2}\leq c_{1}^{2}\log^{2}n$ when $w=c_{2}\log N$ with $c_{2}>0$ sufficiently large. By Markov’s inequality, with probability at least $1-1/(c_{1}^{2}\log^{2}n)\geq 1-1/\log^{2}n$, there is no 3-tree of size larger than $c_{2}\log N\leq c_{2}\log(c_{1}\log n)$. This implies that with probability at least $1-1/\log^{2}n$, there is no connected component of size larger than than $c_{2}(dk)^{3}\log(c_{1}\log n)\leq c_{1}c_{2}(dk)^{3}\log\log n$ in $\widetilde{H}_{1}$. ∎
Lemma D.3.
@@
$\mathbb{P}[A_{e}]\leq 2\binom{(1-\beta_{2})k}{\leq\alpha k}2^{-(1-\beta_{2})k}\leq 2^{\left(h\left(\frac{\alpha}{1-\beta_{2}}\right)-1\right)(1-\beta_{2})k+1}.$
@@
$e\cdot 2^{\left(h\left(\frac{\alpha}{1-\beta_{2}}\right)-1\right)(1-\beta_{2})k+1}\cdot(kd+1)<1,$
@@
  • For every $w\in V$ that shares a clause with $v$, if $a_{w}<a_{v}$, then we add $w$ to $\mathcal{T}_{v}$ as a child of $v$.
@@
Lemma D.4 ([1, Theorem 3.1]).
Fix $c>1$. For every $k,d\geq 1$, there exists $c_{3}>0$ such that for all sufficiently large $n$, $\mathbb{P}[|\mathcal{T}_{v}|>c_{3}\log^{kd+1}(n)]<n^{-2c}$.
In Phase 1, whether $v$ is colored or troubled-1 completely depends on its descendants in $\mathcal{T}_{v}$. Thus, to compute the assignment of $v$ in Phase 1, it suffices to run Phase 1 on all vertices in $\mathcal{T}_{v}$. By Lemma D.4, with probability $1-n^{-2c}$, this step takes $\mathrm{polylog}(n)$ time.
If $v$ gets colored in Phase 1, then we return the color directly. Otherwise, if $v$ becomes troubled-1 after Phase 1, then we compute the connected component of $H$ containing $v$ after Phase 1. Recall from Lemma D.1 that with probability $1-n^{-2c}$, all connected components after Phase 1 has size at most $c_{1}\log n$, so this step takes another $\mathrm{polylog}(n)$ time.
Let $H_{v}$ denote this connected component. By Lemma D.2, if we perform $O(\log n/\log\log n)$ parallel instances of Phase 2 on $H_{v}$, then with probability $1-n^{-2c}$, in one of them $H_{v}$ is further reduced to connected components each of size $c_{2}\log\log n$. We then perform Phase 3 to find a valid coloring on the component that contains $v$, which by Lemma D.3 is guaranteed to succeed and finishes in $2^{O(\log\log n)}=\mathrm{polylog}(n)$ time.
@@
Appendix E Proof of Lemma 5.2
We first check conditions in Theorem 2.10. For sufficiently large $k$, we have:
  • $0.7784\cdot\frac{1}{75}<2(1-0.96)<1-$ 0.778,
  • $16k^{4}d^{5}\leq 16k^{4}2^{k/80}\leq 2^{0.014k}\leq 2^{(\beta_{1}-h(1-\beta_{1}))k}$,
  • $16k^{4}d^{5}\leq 16k^{4}2^{k/80}\leq 2^{0.03k}\leq 2^{(\beta_{2}-\beta_{1})k-h\left(\frac{\beta_{2}-\beta_{1}}{1-\beta_{1}}\right)(1-\beta_{1})k}$,
  • $2e(kd+1)<2e(k2^{k/400}+1)\leq 2^{0.003k}\leq 2^{\left(1-h\left(\frac{\alpha}{1-\beta_{2}}\right)\right)(1-\beta_{2})k}$.
Moreover, letting $f(\delta):=(\beta_{2}-\delta)-h\left(\frac{\beta_{2}-\delta}{1-\delta}\right)(1-\delta)$, we have
$f^{\prime}(\delta)$ $=-\frac{\left(1-\frac{0.96\,-\delta}{1-\delta}\right)\log\left(1-\frac{0.96\,-\delta}{1-\delta}\right)}{\log(2)}-\frac{(0.96\,-\delta)\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-\delta)\log(2)}$
$\qquad-(1-\delta)\bigg{(}-\frac{\frac{0.96\,-\delta}{(1-\delta)^{2}}-\frac{1}{1-\delta}}{\log(2)}-\frac{\left(\frac{1}{1-\delta}-\frac{0.96\,-x}{(1-x)^{2}}\right)\log\left(1-\frac{0.96\,-\delta}{1-\delta}\right)}{\log(2)}$
$\qquad\qquad+\frac{\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-x)\log(2)}-\frac{(0.96\,-x)\log\left(\frac{0.96\,-\delta}{1-\delta}\right)}{(1-\delta)^{2}\log(2)}-\frac{\frac{1}{1-\delta}-\frac{0.96\,-\delta}{(1-\delta)^{2}}}{\log(2)}\bigg{)}-1$
$<0$
for $0.7780\leq\delta\leq\beta_{1}=$ 0.778.
We now check conditions in 5.1. For sufficiently large $k$, we have:
  • $k\cdot 2^{-\alpha k}\cdot(dk)^{5}\cdot 4\leq k^{6}2^{k/80-k/75}\cdot 4\leq\frac{1}{150e^{3}}$,
  • $\theta=1-\frac{1}{2}\exp\left(\frac{2edk}{2^{\alpha k}}\right)\geq 1-\frac{1}{2}\exp(2ek2^{k/400}-2^{k/75})\geq 0.4$,
  • $236ed^{3}k^{4}\cdot 0.6^{\alpha k}=36ek^{4}(2^{3/400}\cdot 0.6^{1/75})^{k}\leq 1/$ 2,
  • $2^{-\frac{1}{48dk^{4}}}\cdot e^{\frac{2d^{2}/\alpha}{2^{\alpha k}}}\leq 2^{-\frac{1}{48k^{4}2^{k/400}}}\cdot e^{\frac{150\cdot 2^{k/200}}{2^{k/75}}}=2^{\log_{2}\left(-\frac{1}{48k^{4}2^{k/400}}+\frac{150\log_{2}e\cdot 2^{k/200}}{2^{k/75}}\right)}\leq 0.9$,
  • $d\leq 2^{k/400}\leq 2^{\alpha k/4}.$